# Gravitational wave probes of particle dark matter: a review

### v3

@@ II.1 Generic features of ultralight dark matter signal
If it exists, cold, ultralight DM could interact with standard-model particles in model-dependent ways. Such interactions would cause macroscopic differences in the materials, and, depending on the type of DM, would lead to an observable signal in different components of GW interferometers. While the physics behind each type of DM is different, there are some characteristics that are model-independent. First, the number of ultralight DM particles in a given region in space is gigantic, and can be calculated by attributing the DM energy density $\rho_{\text{DM}}$ to result from ultralight DM:
$N_{0}$ $=$ $\lambda_{\rm C}^{3}\frac{\rho_{\text{DM}}}{m_{\rm DM}c^{2}}=\left(\frac{2\pi\hbar}{m_{\rm DM}v_{0}}\right)^{3}\frac{\rho_{\text{DM}}}{{m_{\rm DM}c^{2}}},$ (1)
@@
where $v_{0}\simeq 220$ km/s is the virial velocity Smith et al. (2007), $\hbar$ is Planck’s reduced constant, $c$ is the speed of light, $m_{\rm DM}$ is the mass of the ultralight dark matter particle, $\lambda_{\rm C}$ is the Compton wavelength of DM, and $\rho_{\text{DM}}=0.3\,\mathrm{GeV}/$m3 is the DM energy density de Salas and Widmark (2021) (unless otherwise noted). The large occupation number $N_{0}$ implies that ultralight DM can be treated as a coherent region, where within that region, DM can be approximated as a plane wave that oscillates at a fixed frequency $f_{0}$, which is given by the DM mass:
$f_{0}=\frac{m_{\rm DM}c^{2}}{2\pi\hbar}\simeq 241\text{ Hz}\left(\frac{m_{\rm DM}}{10^{-12}\text{ eV}/c^{2}}\right),$ (2)
@@
In other words, the observable effect is that the interferometer components will forever oscillate at a fixed frequency since they always sit in the DM field Hall et al. (2018). The DM wave also has a finite coherence time $T_{\text{coh}}$ that relates to the virial velocity of DM:
$T_{\text{coh}}=\frac{4\pi\hbar}{m_{\rm DM}v_{0}^{2}}=1.4\times 10^{4}\text{ s}\left(\frac{10^{-12}\text{ eV}}{m_{\rm DM}}\right),$ (3)
If we observe for a duration longer than the coherence time, the signal will no longer be monochromatic but have its power stochastically distributed at frequencies slightly higher than the characteristic frequency of the DM particle:
$\Delta f_{v}=\frac{1}{2}\left(\frac{v_{0}}{c}\right)^{2}f_{0}\approx 2.94\times 10^{-7}f_{0}.$ (4)
This frequency shift arises because in a given coherent region, the number of DM particles is huge, and all are traveling at different velocities that are Maxwell-Boltzmann distributed around the virial velocity. Furthermore, the DM signal has an associated coherence length $L_{\text{coh}}$:
$L_{\text{coh}}=\frac{2\pi\hbar}{m_{\rm DM}v_{0}}=1.6\times 10^{9}\text{ m}\left(\frac{10^{-12}\text{ eV}/c^{2}}{m_{\rm DM}}\right).$ (5)
If observing the interaction of DM with standard-model model particles in detectors separated by distances less than $L_{\text{coh}}$, which is the case for ground-based detectors, we expect to observe a correlated signal, again at an almost fixed frequency. In Fig. 2, we show an example of what a simulated vector DM signal interacting with one of the LIGO detectors looks like.
@@
Figure 2: Taken from Miller et al. (2021). We show the strain time series $h(t)$ (left) and the Fourier transform of it (right) of a simulated dark photon DM signal. The structure in the amplitude spectrum density in the right-hand plot arises because we have taken the length of the Fourier transform to exceed that of the coherence time of the signal, and the signal is a superposition of $1000$ dark photons traveling with distinct Maxwell-Boltzmann-distributed velocities, which cause small frequency deviations away from the minimal frequency $f_{0}=740.436$ Hz ($m_{A}=3.06\times 10^{-12}$ eV$/c^{2}$). This plot shows why it is important to choose $T_{\text{FFT}}\sim T_{\text{coh}}$ for the analysis of DM signals. Though we have simulated a dark photon signal, the power spectrum will appear similarly for the other kinds of DM interaction signals. Here, $T_{\text{obs}}\sim 10^{5}$ s, meaning the frequency resolution is $\delta f=10^{-5}$ Hz. The coherence time and length of this signal are: $T_{\text{coh}}=4595$ s and $L_{\text{coh}}=5.28\times 10^{8}$ m; the coupling strength is $\epsilon=3\times 10^{-21}$. The signal is actually simulated for $\sim 233$ days, though we only show the first day of its time evolution.
@@
We will now describe different ways in which DM could interact with particles in the interferometers. We will assume that we are working within $T_{\text{coh}}$, in order to avoid technical difficulties from working with particles of slightly different frequencies. However, we note that in practice, this consideration is necessary when modelling the signal and assessing the sensitivity of GW interferometers to DM/standard-model interactions.
@@ II.2 Axions
Axions could couple to photons, given by the following Chern-Simons interaction Lagrangian:
@@
[…]
where $a(t)$ is the axion field, $g_{\text{a}\gamma}$ is the axion-photon coupling constant, $F_{\mu\nu}$ is field strength of electromagnetic field, $\tilde{F}^{\mu\nu}\equiv\epsilon^{\mu\nu\rho\sigma}F_{\rho\sigma}/2$ is its Hodge dual with the Levi-Civita anti-symmetric tensor $\epsilon^{\mu\nu\rho\sigma}$, the dot denotes the time derivative, and $A_{\mu}$ is the electromagnetic vector potential. $F_{\mu\nu}\equiv\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$. All spatial derivatives here are being neglected.
@@
where $\bm{k}$ is the wave number vector, and “L” and “R” denote the left- and right-hand polarizations, and $\bm{x}$ is a spatial point at which the DM field is measured. By completely fixing gauge degrees of freedom ($\partial_{i}A_{i}=0$ and $A_{0}=0$, respectively), and noting how the polarization vectors $e^{\lambda}_{i}(\hat{\bm{k}})$ in different directions relate to each other Nagano et al. (2019), the equations of motion for $A_{\text{L/R}}(t,\bm{k})$ can be derived:
@@
$a(t)=a_{0}\cos(m_{\rm DM}t)=\left(\frac{\hbar\sqrt{2\rho_{\text{DM}}}}{m_{\rm DM}c}\right)\cos(m_{\rm DM}t)$ (10)
@@
For our purposes, in the presence of an axion field, the phase velocities of circularly polarized photons in the interferometers are modulated Nagano et al. (2019). This signal can only be seen if additional but simplistic optical components are added to measure the optical path difference between two orthogonal polarizations of light ($p$- and $s$-) Nagano et al. (2021). In practice, GW detectors work with linearly polarized light ($p$-) that will be modulated by axions, making $s$- polarized light. Furthermore, such deviations in polarizations can also be measured in pulsar timing arrays Xue et al. (2024); Porayko et al. (2024).
@@ II.3 Dilaton scalar dark matter
Dilaton-like DM Preskill et al. (1983); Abbott and Sikivie (1983); Dine and Fischler (1983); Cho and Keum (1998); Cho and Kim (2009); Arvanitaki et al. (2015a); Stadnik and Flambaum (2015a, b, 2016) could cause the rest mass of the electron and other physical constants to oscillate with time. In other words, atoms’ Bohr radii of different interferometer components would be modulated at the dilaton frequency Grote and Stadnik (2019), and would induce changes in the size and index of refraction of the beam splitter Vermeulen et al. (2021) and the arm mirrors Göttel et al. (2024). Since beam-splitter size oscillations would asymmetrically affect both rays, DM effects would lead to a differential phase shift (i.e. a measurable signal). In fact, at first the beam splitter effect was assumed to be the only one that mattered, but recently, the finite differences between the thicknesses of the mirrors in each arm of the interferometer lead to non-negligible differences in the path lengths of the light in the interferometer Göttel et al. (2024).
Similarly to the axion, the scalar ultralight dark-matter field $\phi$ can be written as Arvanitaki et al. (2015a); Derevianko (2018); Vermeulen et al. (2021):
$\phi(t,\vec{r})=\left(\frac{\hbar\sqrt{2\rho_{\rm{DM}}}}{m_{\rm DM}c}\right)\cos\left(\omega t-\vec{k}\cdot\vec{r}+\Psi\right),$ (13)
where $\vec{r}$ is a position vector, $\omega=(m_{\rm DM}c^{2})/\hbar$ is the angular Compton frequency, $\vec{k}=(m_{\rm DM}\vec{v}_{\text{obs}})/\hbar$ is the wave vector,$\Psi$ is a phase factor, and $\vec{v}_{\text{obs}}$ is the velocity of the DM relative to the observer.
The Lagrangian $\mathcal{L}_{\mathrm{int}}$ for this scalar field $\phi$ is Vermeulen et al. (2021):
@@
For comparison, we write a part of the standard model Lagrangian as:
$\mathcal{L}_{\textrm{SM}}\supset-\frac{F_{\mu\nu}F^{\mu\nu}}{4}-m_{e}\bar{\psi}_{e}\psi_{e}-eJ_{\mu}A^{\mu}$ (15)
and we can see that the rest mass of the electron effectively changes in the presence of DM:
@@ II.3.1 Changes to size and indices of refraction of solids
The changes in fundamental constants will lead to a change in length of, or strain on, solid bodies. Noting that the length $l$ of an object is roughly $l\sim N_{A}a_{B}$, where $N_{A}$ is the number of atoms in the body and $a_{B}=(m_{e}\alpha)^{-1}$ is the Bohr radius, we can obtain the strain on a solid body in general:
$\delta l=-N_{A}(m_{e}\alpha)^{-2}(m_{e}\delta\alpha+\alpha\delta m_{e})$ (20)
$\frac{\delta l}{l}=-\left(\frac{\delta\alpha(t)}{\alpha}+\frac{\delta m_{e}(t)}{m_{e}}\right)$ (21)
@@
Then, the differential displacement $\delta(L_{x}-L_{y})$ on the beam splitter in GW interferometers can be derived Grote and Stadnik (2019). Along the $x$-arm, assuming that light is incident on the beam splitter at 45 degrees, a reflectivity of the mirror of 50%, the change in the $x$-arm length, $\delta L_{x}$, is calculated as:
@@
where $w$ is the thickness of the mirrors. Along the $y$-arm, the change in arm length is:
@@
$\approx\left(\frac{1}{\Lambda_{\gamma}}+\frac{1}{\Lambda_{e}}\right)\left(\frac{n\,l\,\hbar\,\sqrt{2\,\rho_{\mathrm{DM}}}}{m_{\rm DM}\,c}\right)\cos\left(\omega t\right),$ (28)
where $n$ and $l$ are the index of refraction and length of the beam splitter, respectively. We only consider the $\delta l$ contribution, not one from $\delta n$, because it is much larger than than the $\delta n$ one.
Additionally, dilatons could be visible in a different part of the interferometers, that is, in the reference cavity. Laser light locked to an optical cavity made out of a solid material will have its frequency modulated due to the presence of a dilaton DM field, compared to light locked to a free-space suspended cavity, whose frequency would remain stable Derevianko (2018); Geraci et al. (2019); Hall and Aggarwal (2022). The strain amplitude of such a signal can be written as:
${h}_{\text{DM}}(t)=\left[\frac{\hbar}{m_{\rm DM}c^{3}}\sqrt{8\pi\rho_{\text{DM}}G}(d_{e}+d_{m_{e}})\right]\cos\omega t$ (29)

### v4

@@ II.1 Generic features of ultralight dark matter signals
Cold, ultralight DM could interact with standard-model particles in many model-dependent ways. Such interactions would cause macroscopic differences in the materials, and, depending on the type of DM, would lead to an observable signal in different components of GW interferometers. While the physics behind each type of DM is different, there are some characteristics that are model-independent. First, the number of ultralight DM particles in a given region in space is gigantic, and can be calculated by attributing the DM energy density $\rho_{\text{DM}}$ to result from ultralight DM:
$N_{0}$ $=$ $\lambda_{\rm C}^{3}\frac{\rho_{\text{DM}}}{m_{\rm DM}}=\left(\frac{2\pi}{m_{\rm DM}v_{0}}\right)^{3}\frac{\rho_{\text{DM}}}{{m_{\rm DM}}},$ (1)
@@
where $v_{0}\simeq 220$ km/s is the virial velocity Smith et al. (2007), $m_{\rm DM}$ is the mass of the ultralight dark matter particle, $\lambda_{\rm C}$ is the Compton wavelength of DM, and $\rho_{\text{DM}}=0.3\,\mathrm{GeV}/$cm3 is the DM energy density de Salas and Widmark (2021). The large occupation number $N_{0}$ implies that ultralight DM can be approximated as a plane wave that oscillates at a fixed frequency $f_{0}$, which is given by the DM mass:
$f_{0}=\frac{m_{\rm DM}}{2\pi}\simeq 241\text{ Hz}\left(\frac{m_{\rm DM}}{10^{-12}\text{ eV}}\right),$ (2)
@@
In other words, the observable effect is that the interferometer components will forever oscillate at a fixed frequency since they always sit in the DM field Pierce et al. (2018).
Despite the fixed frequency implied by Eq. 2, we note that in reality, the DM field is not a perfectly monochromatic wave because the constituent particles have a finite Maxwell-Boltzmann velocity dispersion set by the virial velocity $v_{0}$ of the galactic halo. This spread in velocities causes a small but finite spread in the oscillation frequency, which in turn limits the coherence of the field. As a result, the DM wave can be treated as coherent only over a characteristic time scale $T_{\text{coh}}$, after which the accumulated phase dispersion between different velocity components causes the signal to lose coherence Carney et al. (2021). The wave coherent time $T_{\text{coh}}$ is given by Carney et al. (2021):
$T_{\text{coh}}=\frac{4\pi}{m_{\rm DM}v_{0}^{2}}=1.4\times 10^{4}\text{ s}\left(\frac{10^{-12}\text{ eV}}{m_{\rm DM}}\right),$ (3)
Practically speaking, we can observe the interaction of DM with ground-based GW interferometers for as long as the detectors are on. Since observing runs of LIGO, Virgo and KAGRA span around one year, we will always be in a regime in which the signal will not be monochromatic and thus have its power stochastically distributed at frequencies slightly higher than the characteristic frequency of the DM particle. The following expression sets the range of frequencies for which this happens:
$\Delta f_{v}=\frac{1}{2}v_{0}^{2}f_{0}\approx 2.94\times 10^{-7}f_{0}.$ (4)
Additionally, a finite coherence time also implies a finite coherence length $L_{\text{coh}}$:
$L_{\text{coh}}=\frac{2\pi}{m_{\rm DM}v_{0}}=1.6\times 10^{9}\text{ m}\left(\frac{10^{-12}\text{ eV}}{m_{\rm DM}}\right).$ (5)
This length scale characterizes the maximum separation over which two detectors can still observe the DM-induced oscillations with the same phase. Since the arm lengths and baselines of current ground-based GW interferometers are much smaller than $L_{\text{coh}}$, we expect them to record correlated signals at the same frequencies. However, $L_{\text{coh}}$ increases with decreasing $m_{\rm DM}$, which will not be a problem for spaced-based GW interferometers ($L_{\text{coh}}$ will still exceed arm separation and $T_{\text{coh}}$ will greatly exceed the observation time), but does require careful treatment when using pulsar timing arrays, in which $L_{\text{coh}}$ is of the same order as the separation between pulsars. This will be discussed further in Section III.3.
The finite $L_{\text{coh}}$ and $T_{\text{coh}}$ naturally imply that cross-correlation and other CW methods are natural choices to look for this kind of DM, which will be detailed in Section III.1. Additionally, we show a concrete example of a potential ultralight dark-matter signal in Fig. 3. Here, simulated vector DM particles are interacting with one of the LIGO detectors, which produces a stochastic, narrow-band signal in the frequency domain.
@@
Figure 3: Taken from Miller et al. (2021a). We show the strain time series $h(t)$ (left) and the Fourier transform of it (right) of a simulated dark photon DM signal. The structure in the amplitude spectrum density in the right-hand plot arises because we have taken the length of the Fourier transform to exceed that of the coherence time of the signal. This dark photon DM signal is a superposition of $1000$ dark photons traveling with distinct Maxwell-Boltzmann-distributed velocities, which cause small frequency deviations away from the minimal frequency $f_{0}=740.436$ Hz ($m_{A}=3.06\times 10^{-12}$ eV). This plot shows why it is important to choose $T_{\text{FFT}}\sim T_{\text{coh}}$ for the analysis of DM signals. Though we have simulated a dark photon signal, the power spectrum will appear similarly for the other kinds of DM interaction signals. Here, $T_{\text{obs}}\sim 10^{5}$ s, meaning the frequency resolution is $\delta f=10^{-5}$ Hz. The coherence time and length of this signal are: $T_{\text{coh}}=4595$ s and $L_{\text{coh}}=5.28\times 10^{8}$ m; the coupling strength is $\epsilon=3\times 10^{-21}$. The signal is actually simulated for $\sim 233$ days, though we only show the first day of its time evolution.
@@
We will now describe different ways in which DM could interact with particles in the interferometers. We will assume that we are working within $T_{\text{coh}}$, in order to avoid technical difficulties from working with particles of slightly different frequencies. However, we note that in practice, this consideration is necessary when assessing the sensitivity of GW interferometers to DM/standard-model interactions Pierce et al. (2018); Miller et al. (2021a); Vermeulen et al. (2021); Göttel et al. (2024).
@@ II.2 Axions
Axions were originally proposed as a solution to the strong CP problem in QCD Peccei and Quinn (1977a, b), and more generally, axion-like particles arise in many extensions of the standard model. If sufficiently light and weakly interacting, they can also constitute a viable component of DM. A key feature of axions is their ability to couple to standard-model fields, in particular photons, through a Chern–Simons interaction. This coupling is described by the Lagrangian Chern and Simons (1974):
@@
$=g_{\text{a}\gamma}\dot{a}(t)\epsilon_{ijk}A_{i}\partial_{j}A_{k},$ (6)
where $a(t)$ is the axion field, $g_{\text{a}\gamma}$ is the axion-photon coupling constant, $F_{\mu\nu}$ is field strength of electromagnetic field, $\tilde{F}^{\mu\nu}\equiv\epsilon^{\mu\nu\rho\sigma}F_{\rho\sigma}/2$ is its Hodge dual with the Levi-Civita anti-symmetric tensor $\epsilon^{\mu\nu\rho\sigma}$, the dot denotes the time derivative, $\epsilon_{ijk}$ is Levi-Civita tensor, and $A_{\mu}$ is the electromagnetic vector potential. $F_{\mu\nu}\equiv\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$. Because the coherence length of the axion is much larger than typical experimental setups, the axion field can be treated as homogeneous; thus, all spatial derivatives of the axion field here are neglected.
@@
where $\bm{k}$ is the wave number vector, $k\equiv|\bm{k}|$, and “L” and “R” denote the left- and right-hand polarizations, and $\bm{x}$ is a spatial point at which the DM field is measured. By completely fixing gauge degrees of freedom in which the divergence of the spatial part of the field is zero and its temporal component vanishes ($\partial_{i}A_{i}=0$ and $A_{0}=0$), and noting that the following conditions are satisfied:
  1. 1.
[…]
  2. 2.
    The two circularly polarized states are orthonormal: $e^{\lambda}_{i}(\hat{\bm{k}})e^{\lambda^{\prime}*}_{i}(\hat{\bm{k}})=\delta^{\lambda\lambda^{\prime}}$
  3. 3.
[…]
the equations of motion for $A_{\text{L/R}}(t,\bm{k})$ can be derived Nagano et al. (2019):
@@
$a(t)=a_{0}\cos(m_{\rm DM}t)=\left(\frac{\sqrt{2\rho_{\text{DM}}}}{m_{\rm DM}}\right)\cos(m_{\rm DM}t)$ (10)
@@
Eq. 11 implies that in the presence of an axion field, the phase velocities of circularly polarized photons in interferometers are modulated. In order to detect this effect in LIGO, Virgo and KAGRA instruments, additional but simplistic optical components are required in order to measure the optical path difference between two orthogonal polarizations of light ($p$- and $s$-polarizations) Nagano et al. (2021). In practice, GW interferometers use linearly polarized light ($p$-polarized), which becomes partially $s$-polarized due to axion-induced modulation. Similar deviations in polarization can also be probed with pulsar timing arrays Xue et al. (2024); Porayko et al. (2025), and experimental constraints on axions from them will be discussed in Section III.3.5. Notably, ground-based interferometers and pulsar timing arrays are sensitive to very different axion mass ranges, roughly $10^{-13},10^{-11}]$ eV and $10^{-23},10^{-21}]$ eV, respectively, but can achieve comparable absolute constraints on the axion-photon coupling $g_{\text{a}\gamma}$, which vary strongly with the axion mass. However, no constraints exist yet for this effect from ground-based interferometers.
@@ II.3 Dilaton scalar dark matter
Another class of ultralight dark matter candidates are dilatons, scalar fields that can couple to standard-model particles and induce time-dependent variations in fundamental constants Preskill et al. (1983); Abbott and Sikivie (1983); Dine and Fischler (1983); Cho and Keum (1998); Cho and Kim (2009); Arvanitaki et al. (2015b); Stadnik and Flambaum (2015a, b, 2016). In particular, a dilaton field can cause the electron mass and other physical constants to oscillate at the dilaton frequency. Similarly to the axion, the scalar ultralight dark-matter field $\phi$ can be written as Arvanitaki et al. (2015b); Derevianko (2018); Vermeulen et al. (2021):
$\phi(t,\vec{x})=\left(\frac{\sqrt{2\rho_{\rm{DM}}}}{m_{\rm DM}}\right)\cos\left(m_{\rm DM}t-\vec{k}\cdot\vec{x}+\Psi\right),$ (13)
where $\vec{x}$ is a position vector, $\vec{k}=m_{\rm DM}\vec{v}_{0}$ is the wave vector, and $\Psi$ is a phase factor, and has a corresponding Lagrangian $\mathcal{L}_{\mathrm{int}}$ Vermeulen et al. (2021):
@@
We will now show how the fundamental constants will change in response to the dilaton field. To begin, we write a part of the standard model Lagrangian that will be compared to Eq. 14:
$\mathcal{L}_{\textrm{SM}}\supset-\frac{F_{\mu\nu}F^{\mu\nu}}{4}-m_{e}\bar{\psi}_{e}\psi_{e}-eJ_{\mu}A^{\mu},$ (15)
where $J^{\mu}$ is a current term. From comparing Eq. 14 and Eq. 15 term-wise, we can see that the rest mass of the electron effectively changes in the presence of DM:
@@
Thus, the fundamental constants oscillate in response to the dilaton field. In the following subsections, we will discuss how these variations induce measurable strain in ground-based interferometers and pulsar timing arrays through their effects on solids, neutron stars, and atomic clocks.
@@ II.3.1 Changes to size and indices of refraction of solids
The oscillations in fundamental constants induced by a dilaton field lead to modulations of the Bohr radii of atoms in interferometer components Grote and Stadnik (2019), leading to measurable changes in the size and refractive index of the beam splitter Vermeulen et al. (2021), and the arm mirrors Göttel et al. (2024). Oscillations in the beam splitter size cause photons traveling down each arm to traverse slightly different paths along its surface, producing a differential phase shift. Initially, it was assumed that the beam-splitter effect was the dominant contribution Vermeulen et al. (2021), but more recent work has shown that small differences in the thicknesses of the mirrors in each arm also lead to non-negligible path-length variations in the interferometer Göttel et al. (2024).
We will now calculate the impact of dilatons on GW interferometers. A typical solid length can be estimated as roughly $l\sim N_{A}a_{B}$, where $N_{A}$ is the number of atoms in the body and $a_{B}=(m_{e}\alpha)^{-1}$ is the Bohr radius. Small changes in the electron mass or fine-structure constant then produce a fractional change in length, or strain, given by
$\delta l$ $=-N_{A}(m_{e}\alpha)^{-2}(m_{e}\delta\alpha+\alpha\delta m_{e})$
$\frac{\delta l}{l}$ $=-\left(\frac{\delta\alpha(t)}{\alpha}+\frac{\delta m_{e}(t)}{m_{e}}\right)$
@@
In the context of GW interferometers, this effect translates into a differential displacement of the beam splitter along the two arms. For the $x$-arm, assuming light incidence at 45 degrees and 50% mirror reflectivity, the induced change in arm length, $\delta L_{x}$, can be computed as in Grote and Stadnik (2019):
@@
where $w$ is the thickness of the mirrors, $n$ is the index of the beamsplitter, and $l$ is the length of the beamsplitter. Along the $y$-arm, the change in arm length is Vermeulen et al. (2021):
@@
$\approx\left(\frac{1}{\Lambda_{\gamma}}+\frac{1}{\Lambda_{e}}\right)\left(\frac{n\,l\,\sqrt{2\,\rho_{\mathrm{DM}}}}{m_{\rm DM}\,}\right)\cos\left(m_{\rm DM}t\right),$ (26)
We have only considered the contribution arising from the change in length of the beamsplitter, not from the change in index of refraction, because the former is much larger than the latter Grote and Stadnik (2019).
In addition to dilatons interacting with atoms in the beamsplitter, the reference cavity of the detector can provide a region in which dilatons interact appreciably with light, allowing for potential detection. Laser light locked to an optical cavity made out of a solid material will have its frequency modulated due to the presence of a dilaton DM field, compared to light locked to a free-space suspended cavity, whose frequency would remain stable Derevianko (2018); Geraci et al. (2019); Hall and Aggarwal (2022). The strain amplitude of such a signal, which would be distinct from that in Eq. 26, can be written as:
${h}_{\text{DM}}(t)=\left[\frac{1}{m_{\rm DM}}\sqrt{8\pi\rho_{\text{DM}}G}(d_{e}+d_{m_{e}})\right]\cos m_{\rm DM}t,$ (27)
@@
The effects described in Eq. 26 and Eq. 27 can both produce measurable signals in interferometers. Currently, however, only the spatial strain affecting mirrors and the beamsplitter has motivated dedicated search methods for dilaton dark matter (see Section III.1) and has been constrained using data from GEO600 and LIGO (see Section III.2). A search strategy targeting the temporal modulation of the laser frequency is still under development Hall and Aggarwal (2022).